OpenStudy (anonymous):

Solve the proportion: 16/17 = x/40

OpenStudy (anonymous):

cross multiply ? i think so. are there answer choices? 16/17 = X/40 17x = 640 x = 640/17 = 37.6

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

so 37.6? Thank you, I didn't know how to get the answer! 37.6 is one of the answer choices.

OpenStudy (anonymous):

oh! your welcome :)